three motorcyclist rides around a circular race course starting at the same place and the same time. The first passes the starting point every 12 mins, the second every 18 mins, and the third every 16 mins. After how many mins will all three pass the starting point again at the same time. Explain.
Answer by ankor@dixie-net.com(22740)

three motorcyclist rides around a circular race course starting at the same place
and the same time.
The first passes the starting point every 12 mins, the second every 18 mins, and
the third every 16 mins.
After how many mins will all three pass the starting point again at the same time.
:
Find the least common multiple of 12,18, 16
:
Prime factor
12: 2*2*3
18: 2*3*3
16: 2*2*2*2
:
LCM: 2*2*2*2*3*3 = 144 min all three will pass the starting point at the same time
:
:
#1: 144/12 = 12 laps
#2: 144/18 = 8 laps
#3: 144/16 = 9 laps